$SPAD/src/input zimmbron.input

Paul Zimmermann and Manuel Bronstein

December 3, 2016

Abstract

Contents

2

— * —

)set break resume
)sys rm -f zimmbron.output
)spool zimmbron.output
)set message test on
)set message auto off
)clear all
--S 1 of 143
y := operator y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 1
--S 2 of 143
y' := D(y x, x)
--R
--R
--R         ,
--R   (2)  y (x)
--R
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 143
y'' := D(y', x)
--R
--R
--R         ,,
--R   (3)  y  (x)
--R
--R                                                    Type: Expression(Integer)
--E 3
--S 4 of146
y1 := operator y1
--R
--R
--R   (4)  y1
--R                                                          Type: BasicOperator
--E 4
--S 5 of 143
y2 := operator y2
--R
--R

--R   (5)  y2
--R                                                          Type: BasicOperator
--E 5
--S 6 of 143
y1' := D(y1 t, t)
--R
--R
--R          ,
--R   (6)  y1 (t)
--R
--R                                                    Type: Expression(Integer)
--E 6
--S 7 of 143
y2' := D(y2 t, t)
--R
--R
--R          ,
--R   (7)  y2 (t)
--R
--R                                                    Type: Expression(Integer)
--E 7

———-

0.1 First Order 1

— * —

--S 8 of 143
eq := (x^4 - x^3) * y' + 2 * x^4 * y x = x^3/3 + C
--R
--R
--R                                   3
--R          4    3  ,        4      x  + 3C
--R   (8)  (x  - x )y (x) + 2x y(x)= -------
--R                                     3
--R                                          Type: Equation(Expression(Integer))
--E 8
--S 9 of 143
solve(eq,y,x)
--R
--R
--R                         3     2                     - 2x
--R                       2x  - 3x  + 6C              %e
--R   (9)  [particular= ------------------,basis= [-----------]]

--R                        4      3      2          2
--R                     12x  - 24x  + 12x          x  - 2x + 1
--IType: Union(Record(particular: Expression(Integer),basis: ...
--E 9

———-

0.2 First Order 2

— * —

--S 10 of 143
eq := - y' / 2 + y x = sin x
--R
--R
--R            ,
--R         - y (x) + 2y(x)
--R
--R   (10)  ---------------= sin(x)
--R                2
--R                                          Type: Equation(Expression(Integer))
--E 10
--S 11 of 143
solve(eq,y,x)
--R
--R
--R                      4sin(x) + 2cos(x)           2x
--R   (11)  [particular= -----------------,basis= [%e  ]]
--R                              5
--IType: Union(Record(particular: Expression(Integer),basis: ...
--E 11

———-

0.3 First Order 3

— * —

--S 12 of 143
eq := y' = y x / (y x * log y x + x)
--R
--R
--R          ,            y(x)
--R   (12)  y (x)= -----------------
--R                y(x)log(y(x)) + x

--R                                          Type: Equation(Expression(Integer))
--E 12
--S 13 of 143
solve(eq,y,x)
--R
--R
--R                      2
--R         y(x)log(y(x))  - 2x
--R   (13)  -------------------
--R                2y(x)
--R                                         Type: Union(Expression(Integer),...)
--E 13

———-

0.4 First Order 4

— * —

--S 14 of 143
eq := 2*(y x)*y'**2-2*x*y'-y x=0
--R
--R
--R               ,   2      ,
--R   (14)  2y(x)y (x)  - 2xy (x) - y(x)= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 14
--S 15 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 15

———-

0.5 Bernoulli

— * —

--S 16 of 143
eq := y' + y x = y(x)^3 * sin x
--R
--R
--R          ,                3
--R   (15)  y (x) + y(x)= y(x) sin(x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 16
--S 17 of 143
solve(eq,y,x)
--R
--R
--R                2              2
--R         - 4y(x) sin(x) - 2y(x) cos(x) + 5
--R   (16)  ---------------------------------
--R                          2  2x
--R                     5y(x) %e
--R                                         Type: Union(Expression(Integer),...)
--E 17

———-

0.6 Bernoulli 2

— * —

--S 18 of 143
P := operator P
--R
--R
--R   (17)  P
--R                                                          Type: BasicOperator
--E 18
--S 19 of 143
Q := operator Q
--R
--R
--R   (18)  Q
--R                                                          Type: BasicOperator
--E 19
--S 20 of 143
eq := y' + P x * y x = Q x * y(x)^n
--R
--R

--R          ,                        n
--R   (19)  y (x) + P(x)y(x)= Q(x)y(x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 20
--S 21 of 143
solve(eq,y,x)
--R
--R
--R   (20)  "failed"
--R                                                    Type: Union("failed",...)
--E 21

———-

In the original text this gave the wrong answer. The answer should contain Q. The original text read:

                  1
   (19)  --------------------
                  x
                ++
                |   P(%L)d%L
               ++
         y(x)%e

— * —

--S 22 of 143
eq := y' + P x * y x = Q x * y(x)^2
--R
--R
--R          ,                        2
--R   (21)  y (x) + P(x)y(x)= Q(x)y(x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 22
--S 23 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   Function not supported by Risch d.e.
--R
--R   Continuing to read the file...

--R
--E 23

———-

We cannot solve for n rational — * —

--S 24 of 143
eq := y' + P x * y x = Q x * y(x)^(2/3)
--R
--R
--R          ,                    3+----+2
--R   (22)  y (x) + P(x)y(x)= Q(x)\|y(x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 24
--S 25 of 143
solve(eq,y,x)
--R
--R
--R   (23)  "failed"
--R                                                    Type: Union("failed",...)
--E 25

———-

Again, the previous answer was wrong. It should contain Q. The original result was:

                   1
   (31)  ----------------------
                   x
                 ++
                 |   2P(%L)d%L
                ++
             2
         y(x) %e

— * —

--S 26 of 143
eq := y' + P x * y x = Q x * y(x)^3
--R
--R
--R          ,                        3
--R   (24)  y (x) + P(x)y(x)= Q(x)y(x)
--R

--S 27 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   Function not supported by Risch d.e.
--R
--R   Continuing to read the file...
--R
--E 27

———-

0.7 Clairaut

— * —

--S 28 of 143
eq := (x^2-1)*y'^2-2*x*y(x)*y'+y(x)^2-1=0
--R
--R
--R           2      ,   2           ,          2
--R   (25)  (x  - 1)y (x)  - 2x y(x)y (x) + y(x)  - 1= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 28
--S 29 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 29

———-

0.8 Clairaut2

--S 30 of 143
f := operator f
--R
--R
--R   (26)  f
--R                                                          Type: BasicOperator
--E 30
--S 31 of 143
g := operator g
--R
--R
--R   (27)  g
--R                                                          Type: BasicOperator
--E 31
--S 32 of 143
eq := f(x*y'-y(x))=g(y')
--R
--R
--R             ,                ,
--R   (28)  f(xy (x) - y(x))= g(y (x))
--R
--R                                          Type: Equation(Expression(Integer))
--E 32
--S 33 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseODE: equation has order 0
--R
--R   Continuing to read the file...
--R
--E 33

———-

0.9 Exact First

— * —

--S 34 of 143
eq := y' = (3 * x^2 - y(x)^2 - 7) / (exp y x + 2 * x * y x + 1)

--R
--R
--R                        2     2
--R          ,       - y(x)  + 3x  - 7
--R   (29)  y (x)= --------------------
--R                  y(x)
--R                %e     + 2x y(x) + 1
--R                                          Type: Equation(Expression(Integer))
--E 34
--S 35 of 143
solve(eq,y,x)
--R
--R
--R           y(x)         2           3
--R   (30)  %e     + x y(x)  + y(x) - x  + 7x
--R                                         Type: Union(Expression(Integer),...)
--E 35

———-

0.10 Homogeneous

— * —

--S 36 of 143
eq := y' = (2 * x^3 * y x - y(x)^4) / (x^4 - 2 * x * y(x)^3)
--R
--R
--R                    4     3
--R          ,     y(x)  - 2x y(x)
--R   (31)  y (x)= ---------------
--R                        3    4
--R                 2x y(x)  - x
--R                                          Type: Equation(Expression(Integer))
--E 36
--S 37 of 143
solve(eq,y,x)
--R
--R
--R   (32)  "failed"
--R                                                    Type: Union("failed",...)
--E 37

———-

0.11 Factor

— * —

--S 38 of 143
eq := y'*(y'+y(x))=x*(x+y(x))
--R
--R
--R          ,   2        ,               2
--R   (33)  y (x)  + y(x)y (x)= x y(x) + x
--R
--R                                          Type: Equation(Expression(Integer))
--E 38
--S 39 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 39

———-

0.12 Interchange

— * —

--S 40 of 143
eq := y' = x / (x^2 * y(x)^2 + y(x)^5)
--R
--R
--R          ,            x
--R   (34)  y (x)= ---------------
--R                    5    2    2
--R                y(x)  + x y(x)
--R                                          Type: Equation(Expression(Integer))
--E 40
--S 41 of 143
solve(eq,y,x)
--R
--R
--R                                      3

--R                                 2y(x)
--R                               - ------
--R                 3     2            3
--R         (- 2y(x)  - 2x  - 3)%e
--R   (35)  ------------------------------
--R                        4
--R                                         Type: Union(Expression(Integer),...)
--E 41

———-

0.13 Lagrange

— * —

--S 42 of 143
eq := y(x)=2*x*y'-a*y'^3
--R
--R
--R                    ,   3      ,
--R   (36)  y(x)= - a y (x)  + 2xy (x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 42
--S 43 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 43

———-

0.14 Lagrange 2

— * —

--S 44 of 143
eq := y(x)=2*x*y'-y'^2
--R

--R
--R                  ,   2      ,
--R   (37)  y(x)= - y (x)  + 2xy (x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 44
--S 45 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 45

———-

0.15 Riccati

— * —

--S 46 of 143
eq := y' = exp x * y(x)^2 - y x + exp(-x)
--R
--R
--R          ,         2  x     - x
--R   (38)  y (x)= y(x) %e  + %e    - y(x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 46
--S 47 of 143
solve(eq,y,x)
--R
--R
--R   (39)  "failed"
--R                                                    Type: Union("failed",...)
--E 47

———-

0.16 Riccati 2

--S 48 of 143
eq := y' = y(x)^2 - x * y x + 1
--R
--R
--R          ,         2
--R   (40)  y (x)= y(x)  - x y(x) + 1
--R
--R                                          Type: Equation(Expression(Integer))
--E 48
--S 49 of 143
solve(eq,y,x)
--R
--R
--R                          2
--R                         x
--R                       - --   x
--R                          2 ++       1
--R         (- y(x) + x)%e     |   - ------- d%M  + 1
--R                           ++           2
--R                                      %M
--R                                    - ---
--R                                       2
--R                                  %e
--R   (41)  -----------------------------------------
--R                                     2
--R                                    x
--R                                  - --
--R                                     2
--R                      (y(x) - x)%e
--R                                         Type: Union(Expression(Integer),...)
--E 49

———-

0.17 Separable

— * —

--S 50 of 143
eq := y' = (9 * x^8 + 1) / (y(x)^2 + 1)
--R
--R
--R                   8
--R          ,      9x  + 1
--R   (42)  y (x)= ---------

--R                    2
--R                y(x)  + 1
--R                                          Type: Equation(Expression(Integer))
--E 50
--S 51 of 143
solve(eq,y,x)
--R
--R
--R             3             9
--R         y(x)  + 3y(x) - 3x  - 3x
--R   (43)  ------------------------
--R                     3
--R                                         Type: Union(Expression(Integer),...)
--E 51

———-

0.18 Solvable x

— * —

--S 52 of 143
eq := y(x) = 2*x*y'+y(x)*y'^2
--R
--R
--R                    ,   2      ,
--R   (44)  y(x)= y(x)y (x)  + 2xy (x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 52
--S 53 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 53

———-

0.19 Solvable y

— * —

--S 54 of 143
eq := x = y(x)*y'-x*y'^2
--R
--R
--R                 ,   2        ,
--R   (45)  x= - x y (x)  + y(x)y (x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 54
--S 55 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 55

———-

0.20 Second Order Change of Variable

— * —

--S 56 of 143
eq := y''*(a*x+b)^2+4*y'*(a*x+b)*a+2*y(x)*a^2 = 0
--R
--R
--R           2 2             2  ,,         2          ,        2
--R   (46)  (a x  + 2a b x + b )y  (x) + (4a x + 4a b)y (x) + 2a y(x)= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 56
--S 57 of 143
solve(eq,y,x)
--R
--R
--R                                         x              2a x + b
--R   (47)  [particular= 0,basis= [------------------,------------------]]

--R                                 2 2             2  2 2             2
--R                                a x  + 2a b x + b  a x  + 2a b x + b
--IType: Union(Record(particular: Expression(Integer),basis: ...
--E 57

———-

0.21 Adjoint

— * —

--S 58 of 143
eq := (x^2-x)*y'' + (2*x^2+4*x-3)*y' + 8*x*y(x) = 1
--R
--R
--R           2      ,,         2           ,
--R   (48)  (x  - x)y  (x) + (2x  + 4x - 3)y (x) + 8x y(x)= 1
--R
--R                                          Type: Equation(Expression(Integer))
--E 58
--S 59 of 143
solve(eq,y,x)
--R
--R
--R                          3     2                                   - 2x
--R                        2x  - 3x  + 53                 1          %e
--R   (49)  [particular= ------------------,basis= [-------------,-----------]]
--R                         4      3      2          4     3    2  2
--R                      12x  - 24x  + 12x          x  - 2x  + x  x  - 2x + 1
--IType: Union(Record(particular: Expression(Integer),basis: ...
--E 59

———-

0.22 Second Order

— * —

--S 60 of 143
eq := (x^2-x)*y'' + (1-2*x^2)*y' + (4*x-2)*y(x) = 0
--R
--R
--R           2      ,,           2      ,
--R   (50)  (x  - x)y  (x) + (- 2x  + 1)y (x) + (4x - 2)y(x)= 0
--R

--S 61 of 143
solve(eq,y,x)
--R
--R
--R                                 2   2x
--R   (51)  [particular= 0,basis= [x ,%e  ]]
--IType: Union(Record(particular: Expression(Integer),basis: ...
--E 61

———-

0.23 Autonomous

— * —

--S 62 of 143
eq := y''-y' = 2*y(x)*y'
--R
--R
--R          ,,       ,           ,
--R   (52)  y  (x) - y (x)= 2y(x)y (x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 62
--S 63 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getfreelincoeff: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 63

———-

0.24 Autonomous 2

— * —

--S 64 of 143
eq := y''/y(x)-y'^2/y(x)^2-1+1/y(x)^3 = 0
--R
--R
--R             2 ,,           ,   2       3
--R         y(x) y  (x) - y(x)y (x)  - y(x)  + 1
--R
--R   (53)  ------------------------------------= 0
--R                             3
--R                         y(x)
--R                                          Type: Equation(Expression(Integer))
--E 64
--S 65 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseLODE: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 65

———-

0.25 y missing

— * —

--S 66 of 143
eq := y'' + 2 * x * y' = 2 * x
--R
--R
--R          ,,         ,
--R   (54)  y  (x) + 2xy (x)= 2x
--R
--R                                          Type: Equation(Expression(Integer))
--E 66
--S 67 of 143
solve(eq,y,x)
--R
--R
--R   (55)  [particular= x,basis= [1,erf(x)]]
--IType: Union(Record(particular: Expression(Integer),basis: ...
--E 67

———-

0.26 diff

— * —

--S 68 of 143
eq := 2*y(x)*y''-y'^2 = (y'-x*y'')^2/3
--R
--R
--R                                2 ,,   2      ,    ,,       ,   2
--R                               x y  (x)  - 2xy (x)y  (x) + y (x)
--R               ,,       ,   2
--R   (56)  2y(x)y  (x) - y (x) = ----------------------------------
--R                                                3
--R                                          Type: Equation(Expression(Integer))
--E 68
--S 69 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 69

———-

0.27 Equidimensional x

— * —

--S 70 of 143
eq := x*y'' = 2*y(x)*y'
--R
--R
--R           ,,           ,
--R   (57)  xy  (x)= 2y(x)y (x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 70

--S 71 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getfreelincoeff: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 71

———-

0.28 Equidimensional y

— * —

--S 72 of 143
eq := (1-x)*(y(x)*y''-y'^2)+x^2*y(x)^2 = 0
--R
--R
--R                       ,,              ,   2    2    2
--R   (58)  (- x + 1)y(x)y  (x) + (x - 1)y (x)  + x y(x) = 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 72
--S 73 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseLODE: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 73

———-

0.29 Exact 2nd Order

— * —

--S 74 of 143
eq := x*y(x)*y''+x*y'^2+y(x)*y'=0
--R
--R
--R                ,,         ,   2        ,
--R   (59)  x y(x)y  (x) + x y (x)  + y(x)y (x)= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 74
--S 75 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseLODE: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 75

———-

0.30 Factoring

— * —

--S 76 of 143
eq := y''^2-2*y'*y''+2*y(x)*y'-y(x)^2=0
--R
--R
--R          ,,   2     ,    ,,            ,          2
--R   (60)  y  (x)  - 2y (x)y  (x) + 2y(x)y (x) - y(x) = 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 76
--S 77 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R

--E 77

———-

0.31 Liouvillian

The original document gave a different answer.

                                                           1
                                               1           --
                                             - -   x       %L
                                               x ++      %e
                                      1    %e    |   ------------- d%L
                                    - -         ++    +----------+
                                      x               |  4      3
                                  %e                 \|%L  - 2%L
   (65)  [particular= 0,basis= [----------,----------------------------]]
                                 +-------+           +-------+
                                 | 2                 | 2
                                \|x  - 2x           \|x  - 2x

— * —

--S 78 of 143
eq := (x^3/2-x^2)*y'' + (2*x^2-3*x+1)*y' + (x-1)*y(x) = 0
--R
--R
--R           3     2  ,,         2           ,
--R         (x  - 2x )y  (x) + (4x  - 6x + 2)y (x) + (2x - 2)y(x)
--R
--R   (61)  -----------------------------------------------------= 0
--R                                   2
--R                                          Type: Equation(Expression(Integer))
--E 78
--S 79 of 143
solve(eq,y,x)
--R
--R
--R   (62)
--R   [particular= 0,
--R                1               1                  1
--R              - - +-------+   - - +-------+   x   -- +----------+
--R                x |   1         x |   1     ++    %M |     1
--R    basis= [%e    |------- ,%e    |-------  |   %e   |---------- d%M ]]
--R                  | 2             | 2      ++        |  4      3
--R                 \|x  - 2x       \|x  - 2x          \|%M  - 2%M

--IType: Union(Record(particular: Expression(Integer),...
--E 79

———-

0.32 Reduction

— * —

--S 80 of 143
eq := y'' - 2 * x * y' + 2 * y x = 3
--R
--R
--R          ,,         ,
--R   (63)  y  (x) - 2xy (x) + 2y(x)= 3
--R
--R                                          Type: Equation(Expression(Integer))
--E 80
--S 81 of 143
solve(eq,y,x)
--R
--R
--R                                            2
--R                                      x   %M
--R                      3             ++  %e
--R   (64)  [particular= -,basis= [x,x |   ----- d%M ]]
--R                      2            ++      2
--R                                         %M
--IType: Union(Record(particular: Expression(Integer),...
--E 81

———-

0.33 Integer Factors

— * —

--S 82 of 143
eq := sqrt(x) * y'' + 2 * x * y' + 3 * y x = 0
--R
--R
--R          +-+ ,,         ,
--R   (65)  \|x y  (x) + 2xy (x) + 3y(x)= 0
--R
--R                                          Type: Equation(Expression(Integer))

--E 82
--S 83 of 143
solve(eq,y,x)
--R
--R
--R   (66)  [particular= 0,basis= []]
--IType: Union(Record(particular: Expression(Integer),...
--E 83

———-

0.34 Scale Invariant

— * —

--S 84 of 143
eq := x^2*y''+3*x*y'+2*y(x)=1/y(x)^3/x^4
--R
--R
--R          2 ,,         ,                1
--R   (67)  x y  (x) + 3xy (x) + 2y(x)= -------
--R                                      4    3
--R                                     x y(x)
--R                                          Type: Equation(Expression(Integer))
--E 84
--S 85 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 85

———-

0.35 undet

— * —

--S 86 of 143

eq := y'' - 2 / x^2 * y x = 7 * x^4 + 3 * x^3
--R
--R
--R          2 ,,
--R         x y  (x) - 2y(x)
--R                             4     3
--R   (68)  ----------------= 7x  + 3x
--R                 2
--R                x
--R                                          Type: Equation(Expression(Integer))
--E 86
--S 97 of 143
solve(eq,y,x)
--R
--R
--R                        7     6     3               3      3
--R                      3x  + 2x  - 7x  + 14         x  - 1 x  + 2
--R   (69)  [particular= --------------------,basis= [------,------]]
--R                               12x                    x      x
--IType: Union(Record(particular: Expression(Integer),...
--E 97

———-

0.36 variation

— * —

--S 88 of 143
eq := y'' + y x = csc x
--R
--R
--R          ,,
--R   (70)  y  (x) + y(x)= csc(x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 88
--S 89 of 143
solve(eq,y,x)
--R
--R
--R   (71)
--R                            sin(x)                     2
--R   [particular= sin(x)log(----------) - sin(x)log(----------) - x cos(x),
--R                          cos(x) + 1              cos(x) + 1
--R    basis= [cos(x),sin(x)]]

———-

0.37 Constant Coefficients

— * —

--S 90 of 143
eq := D(y x,x,7) - 14 * D(y x,x,6) + 80 * D(y x,x,5) - _
       242 * D(y x,x,4) + 419 * D(y x,x,3) - 416 * y'' + _
        220 * y' - 48 * y x = 0
--R
--R
--R   (72)
--R      (vii)         (vi)         (v)          (iv)          ,,,          ,,
--R     y     (x) - 14y    (x) + 80y   (x) - 242y    (x) + 419y   (x) - 416y  (x)
--R
--R   +
--R         ,
--R     220y (x) - 48y(x)
--R
--R     =
--R     0
--R                                          Type: Equation(Expression(Integer))
--E 90
--S 91 of 143
solve(eq,y,x)
--R
--R
--R                                  4x   3x   2x     2x   x     x  2  x
--R   (73)  [particular= 0,basis= [%e  ,%e  ,%e  ,x %e  ,%e ,x %e ,x %e ]]
--IType: Union(Record(particular: Expression(Integer),...
--E 91

———-

0.38 Euler

— * —

--S 92 of 143
eq := D(y x,x,4) - 4/x^2 * y'' + 8/x^3 * y' -8/x^4 * y x = 0
--R

--R
--R          4 (iv)        2 ,,         ,
--R         x y    (x) - 4x y  (x) + 8xy (x) - 8y(x)
--R
--R   (74)  ----------------------------------------= 0
--R                             4
--R                            x
--R                                          Type: Equation(Expression(Integer))
--E 92
--S 93 of 143
solve(eq,y,x)
--R
--R
--R   (75)
--R   [particular= 0,
--R
--R     basis =
--R         5     3     2       5     3      2       5      3     2
--R        x  - 5x  + 5x  - 1  x  + 5x  - 10x  + 4  x  - 10x  + 5x  + 4
--R       [------------------, -------------------, -------------------,
--R                 x                   x                    x
--R         5      3      2
--R        x  - 10x  + 20x  + 4
--R        --------------------]
--R                  x
--R     ]
--IType: Union(Record(particular: Expression(Integer),...
--E 93

———-

0.39 Exact nth

The original document gave a different answer for the particular:

                          2
                         x  - 4                1        x + 1
   (85)  [particular= ------------,basis= [----------,----------,1]]
                        2                   2          2
                      4x  + 4x + 4         x  + x + 1 x  + x + 1

— * —

--S 94 of 143
eq := (1+x+x^2)*D(y x,x,3) + (3+6*x)*y'' + 6*y' = 6*x
--R

--R
--R           2          ,,,               ,,        ,
--R   (76)  (x  + x + 1)y   (x) + (6x + 3)y  (x) + 6y (x)= 6x
--R
--R                                          Type: Equation(Expression(Integer))
--E 94
--S 95 of 143
solve(eq,y,x)
--R
--R
--R                          4
--R                         x  - 4                 1        x + 1
--R   (77)  [particular= ------------,basis= [----------,----------,1]]
--R                        2                   2          2
--R                      4x  + 4x + 4         x  + x + 1 x  + x + 1
--IType: Union(Record(particular: Expression(Integer),...
--E 95

———-

0.40 Circle

— * —

--S 96 of 143
eq := (y'^2+1)*D(y(x),x,3)-3*y'*y''^2 = 0
--R
--R
--R           ,   2      ,,,        ,    ,,   2
--R   (78)  (y (x)  + 1)y   (x) - 3y (x)y  (x) = 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 96
--S 97 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseLODE: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 97

———-

0.41 transfBernoulli

— * —

--S 98 of 143
eq := 3*y''*D(y(x),x,4)-5*D(y(x),x,3)^2 = 0
--R
--R
--R           ,,    (iv)        ,,,   2
--R   (79)  3y  (x)y    (x) - 5y   (x) = 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 98
--S 99 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseLODE: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 99

———-

0.42 Delay

— * —

--S 100 of 143
eq := y'+a*y(x-1) = 0
--R
--R
--R          ,
--R   (80)  y (x) + a y(x - 1)= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 100
--S 101 of 143
solve(eq,y,x)
--R
--R
--R            x
--R          ++

--R   (81)   |   a y(%M - 1)d%M  + y(x)
--R         ++
--R                                         Type: Union(Expression(Integer),...)
--E 101

———-

0.43 Several

— * —

--S 102 of 143
eq := D(y(x,a),x,1) = a*y(x,a)
--R
--R
--R   (82)  y  (x,a)= a y(x,a)
--R          ,1
--R                                          Type: Equation(Expression(Integer))
--E 102
--S 103 of 143
solve(eq,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseODE: equation has order 0
--R
--R   Continuing to read the file...
--R
--E 103

———-

0.44 Nth order

— * —

--S 104 of 143
eq := D(y x,x,4) = sin x
--R
--R
--R          (iv)
--R   (83)  y    (x)= sin(x)
--R
--R                                          Type: Equation(Expression(Integer))

--E 104
--S 105 of 143
ini := [0, 0, 0, 0]
--R
--R
--R   (84)  [0,0,0,0]
--R                                               Type: List(NonNegativeInteger)
--E 105
--S 106 of 143
solve(eq,y,x=0,ini)
--R
--R
--R                    3
--R         6sin(x) + x  - 6x
--R   (85)  -----------------
--R                 6
--R                                         Type: Union(Expression(Integer),...)
--E 106

———-

0.45 BesselJ

— * —

--S 107 of 143
eq := x * y'' + y' + 2 * x * y x = 0
--R
--R
--R           ,,       ,
--R   (86)  xy  (x) + y (x) + 2x y(x)= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 107
--S 108 of 143
solve(eq,y,x=0,[1,0])
--R
--R
--R   (87)  "failed"
--R                                                    Type: Union("failed",...)
--E 108

———-

0.46 Separable

— * —

--S 109 of 143
eq := x*y'^2-y(x)^2+1=0
--R
--R
--R            ,   2       2
--R   (88)  x y (x)  - y(x)  + 1= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 109
--S 110 of 143
solve(eq,y,x=0,[1])
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 110
--S 111 of 143
eq := y' = sqrt((y(x)^2-1)/x)
--R
--R
--R                 +---------+
--R                 |    2
--R          ,      |y(x)  - 1
--R   (89)  y (x)=  |---------
--R                \|    x
--R                                          Type: Equation(Expression(Integer))
--E 111
--S 112 of 143
solve(eq,y,x=0,[1])
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   catdef: division by zero
--R
--R   Continuing to read the file...
--R
--E 112

———-

0.47 ic2

— * —

--S 113 of 143
eq := y''+y(x)*y'^3 = 0
--R
--R
--R          ,,           ,   3
--R   (90)  y  (x) + y(x)y (x) = 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 113
--S 114 of 143
solve(eq,y,x=0,[0,2])
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 114
)clear all

———-

0.48 intcomb

— * —

--S 115 of 143
x := operator x
--R
--R
--R   (1)  x
--R                                                          Type: BasicOperator
--E 115

--S 116 of 143
y := operator y
--R
--R
--R   (2)  y
--R                                                          Type: BasicOperator
--E 116
--S 117 of 143
z := operator z
--R
--R
--R   (3)  z
--R                                                          Type: BasicOperator
--E 117
--S 118 of 143
x' := D(x(t),t)
--R
--R
--R         ,
--R   (4)  x (t)
--R
--R                                                    Type: Expression(Integer)
--E 118
--S 119 of 143
y' := D(y(t),t)
--R
--R
--R         ,
--R   (5)  y (t)
--R
--R                                                    Type: Expression(Integer)
--E 119
--S 120 of 143
z' := D(z(t),t)
--R
--R
--R         ,
--R   (6)  z (t)
--R
--R                                                    Type: Expression(Integer)
--E 120
--S 121 of 143
sys:=[x'=-3*y(t)*z(t),y'=3*x(t)*z(t),z'=-x(t)*y(t)]
--R
--R

--R          ,                  ,                ,
--R   (7)  [x (t)= - 3y(t)z(t),y (t)= 3x(t)z(t),z (t)= - x(t)y(t)]
--R
--R                                    Type: List(Equation(Expression(Integer)))
--E 121
--S 122 of 143
solve(sys,[x,y,z],t)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getfreelincoeff: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 122

———-

0.49 MRiccatti

— * —

--S 123 of 143
a := operator a
--R
--R
--R   (8)  a
--R                                                          Type: BasicOperator
--E 123
--S 124 of 143
b := operator b
--R
--R
--R   (9)  b
--R                                                          Type: BasicOperator
--E 124
--S 125 of 143
eq1 := x'=a(t)*(y(t)^2-x(t)^2)+2*b(t)*x(t)*y(t)+2*c*x(t)
--R
--R
--R          ,             2                           2
--R   (10)  x (t)= a(t)y(t)  + 2b(t)x(t)y(t) - a(t)x(t)  + 2c x(t)
--R
--R                                          Type: Equation(Expression(Integer))

--E 125
--S 126 of 143
eq2 := y'=b(t)*(y(t)^2-x(t)^2)-2*a(t)*x(t)*y(t)+2*c*y(t)
--R
--R
--R          ,             2                                    2
--R   (11)  y (t)= b(t)y(t)  + (- 2a(t)x(t) + 2c)y(t) - b(t)x(t)
--R
--R                                          Type: Equation(Expression(Integer))
--E 126
--S 127 of 143
solve([eq1,eq2],[x,y],t)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 127

———-

0.50 Triangular

— * —

--S 128 of 143
eq1 := x'=x(t)*(1+cos(t)/(2+sin(t)))
--R
--R
--R          ,     x(t)sin(t) + x(t)cos(t) + 2x(t)
--R   (12)  x (t)= -------------------------------
--R                           sin(t) + 2
--R                                          Type: Equation(Expression(Integer))
--E 128
--S 129 of 143
eq2 := y'=x(t)-y(t)
--R
--R
--R          ,
--R   (13)  y (t)= - y(t) + x(t)
--R
--R                                          Type: Equation(Expression(Integer))

--E 129
--S 130 of 143
solve([eq1,eq2],[x,y],t)
--R
--R
--R   (14)  "failed"
--R                                                    Type: Union("failed",...)
--E 130

———-

0.51 Vector

— * —

--S 131 of 143
eq1 := x' = 9*x(t)+2*y(t)
--R
--R
--R          ,
--R   (15)  x (t)= 2y(t) + 9x(t)
--R
--R                                          Type: Equation(Expression(Integer))
--E 131
--S 132 of 143
eq2 := y' = x(t)+8*y(t)
--R
--R
--R          ,
--R   (16)  y (t)= 8y(t) + x(t)
--R
--R                                          Type: Equation(Expression(Integer))
--E 132
--S 133 of 143
solve([eq1,eq2],[x,y],t)
--R
--R
--R                                             10t
--R                                       10t %e        7t     7t
--R   (17)  [particular= [0,0],basis= [[%e   ,-----],[%e  ,- %e  ]]]
--R                                             2
--IType: Union(Record(particular: Vector(Expression(Integer)),...
--E 133

———-

0.52 High Order

— * —

--S 134 of 143
eq1 := x'-x(t)+2*y(t)=0
--R
--R
--R          ,
--R   (18)  x (t) + 2y(t) - x(t)= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 134
--S 135 of 143
eq2 := D(x(t),t,2)-2*y'=2*t-cot(2*t)
--R
--R
--R          ,,        ,
--R   (19)  x  (t) - 2y (t)= - cot(2t) + 2t
--R
--R                                          Type: Equation(Expression(Integer))
--E 135
--S 136 of 143
solve([eq1,eq2],[x,y],t)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   solve: not a first order linear system
--R
--R   Continuing to read the file...
--R
--E 136

———-

0.53 Inhomogeneous

— * —

--S 137 of 143
eq1 := x' = -x(t)/t/(t^2+1)+y(t)/t^2/(t^2+1)+1/t
--R
--R
--R                                 3
--R          ,     y(t) - t x(t) + t  + t

--R   (20)  x (t)= ----------------------
--R                         4    2
--R                        t  + t
--R                                          Type: Equation(Expression(Integer))
--E 137
--S 138 of 143
eq2 := y' = -t^2*x(t)/(t^2+1)+(2*t^2+1)*y(t)/t/(t^2+1)+1
--R
--R
--R                   2             3        3
--R          ,     (2t  + 1)y(t) - t x(t) + t  + t
--R   (21)  y (t)= -------------------------------
--R                              3
--R                             t  + t
--R                                          Type: Equation(Expression(Integer))
--E 138
--S 139 of 143
solve([eq1,eq2],[x,y],t)
--R
--R
--R                                                         1    2
--R   (22)  [particular= [log(t) - 1,t log(t) - t],basis= [[-,- t ],[1,t]]]
--R                                                         t
--IType: Union(Record(particular: Vector(Expression(Integer)),...
--E 139
)clear all

———-

0.54 Bronstein

— * —

--S 140 of 143
s := sqrt(-222)
--R
--R
--R         +-----+
--R   (1)  \|- 222
--R                                                        Type: AlgebraicNumber
--E 140
--S 141 of 143
a0 := 104/25*x^10+(274/25-22/15*s)*x^8+(7754/75-68/15*s)*x^6
--R

--R
--R                        +-----+                  +-----+
--R        104  10   - 110\|- 222  + 822  8   - 340\|- 222  + 7754  6
--R   (2)  --- x   + ------------------- x  + -------------------- x
--R         25                75                       75
--R                                            Type: Polynomial(AlgebraicNumber)
--E 141
--S 142 of 143
a0 := a0+(11248/75-194/15*s)*x^4+(29452/75-296/5*s)*x^2-10952/5-148/3*s
--R
--R
--R   (3)
--R                     +-----+                  +-----+
--R     104  10   - 110\|- 222  + 822  8   - 340\|- 222  + 7754  6
--R     --- x   + ------------------- x  + -------------------- x
--R      25                75                       75
--R   +
--R         +-----+                     +-----+                    +-----+
--R   - 970\|- 222  + 11248  4   - 4440\|- 222  + 29452  2   - 740\|- 222  - 32856
--R   --------------------- x  + ---------------------- x  + ---------------------
--R             75                         75                          15
--R                                            Type: Polynomial(AlgebraicNumber)
--E 142
--S 143 of 143
a2 := x^12+2*x^10+151/3*x^8+296/3*x^6+5920/9*x^4+10952/9*x^2+5476/9
--R
--R
--R         12     10   151  8   296  6   5920  4   10952  2   5476
--R   (4)  x   + 2x   + --- x  + --- x  + ---- x  + ----- x  + ----
--R                      3        3         9         9          9
--R                                          Type: Polynomial(Fraction(Integer))
--E 143

———-

This code loops trying to factor the formula.

— * —

--y := operator y
--y' := D(y x, x)
--y'' := D(y', x)
--eq := a2*y'' - a0*y(x)
--solve(eq,y,x)
--R
--R   WARNING (genufact): No known algorithm to factor
--R      8     43808    6   827062184  4   44060595526144  2   27067141014651136
--R     ?  + --------- ?  - --------- ?  - -------------- ?  - -----------------
--R            +-----+         5625               +-----+           31640625

--R          5\|- 222                      253125\|- 222
--R     , trying square-free.
--R   WARNING (genufact): No known algorithm to factor
--R      8   4292  6   16006348  4   58334426144  2   94037902336
--R     ?  + ---- ?  + -------- ?  + ----------- ?  + -----------
--R           25         1875           421875           140625
--R     , trying square-free.
--R   WARNING (genufact): No known algorithm to factor
--R      8      1776    6   151034  4     108731456   2   902641936
--R     ?  - --------- ?  - ------ ?  + ------------ ?  - ---------
--R            +-----+        625            +-----+        390625
--R          5\|- 222                   9375\|- 222
--R     , trying square-free.
--R   WARNING (genufact): No known algorithm to factor
--R        8      7   3426  6   14364  5   903369  4   277938  3   31725044  2
--R       ?  - 18?  + ---- ?  - ----- ?  + ------ ?  - ------ ?  + -------- ?
--R                    25         25         625         125         15625
--R     +
--R         15594264     3118752
--R       - -------- ? + -------
--R           15625       15625
--R     , trying square-free.
--R

———-

0.55 Moussiaux

This equation crashes Axiom silently and is a major bug.

— * —

--solve(15*D(y(x),x)+24*y(x)^2=7*x^(-8/3),y,x)
)spool
)lisp (bye)

———-

References

[1] Zimmermann, Paul “http://www.loria.fr/ zimmerma/ComputerAlgebra/ode comp.axiom”